# A Survey of Multimodal Large Language Model from A Data-centric Perspective

### v1

@@ 6.2.4. Other Noteworthy Tasks
Harnessing Instruction Tuning data can substantially bolster a model’s proficiency in adhering to instructions and enhancing its performance across specific tasks. Given the pivotal role of both quality and quantity in optimizing model performance, the exploration of this area emerges as a crucial research endeavor. The evaluation of data for large-scale models necessitates metrics that are model-agnostic, effectively capturing the distinct characteristics of the data. As highlighted in the preceding paragraph, there is a pressing need to devise metrics tailored specifically for instruction tuning data.
An alternative pragmatic approach involves employing Large Language Models (LLMs) for data evaluation, leveraging the extensive knowledge accrued during their pre-training phase (Du et al., 2023b; Chen et al., 2023f; Xu et al., 2023b; Liu et al., 2023e; Lu et al., 2023c; Parkar et al., 2024). However, utilizing LLMs for data assessment proves to be cost-effective. Despite the success of GPT-based methods for automated data quality evaluation, such approaches often lack interpretability. Hence, further investigation is warranted to gain a deeper understanding of GPT’s efficacy and to delineate the boundaries of employing GPT for data evaluation.
@@
MLLMs often learn from diverse data types sequentially across various training phases (Li et al., 2022b; Liu et al., 2024b, 2023c). Typically, this sequential training starts with initializing the model using pre-trained weights for each modality, followed by pre-training, instruction tuning, and concluding with reinforcement learning from human feedback. It is critical throughout these stages to prevent catastrophic forgetting, ensuring the model retains its foundational language capabilities while acquiring multimodal skills. Therefore, integrating lifelong learning strategies with MLLMs represents a significant and necessary research direction.

### v2

@@ 6.2. Evaluation Datasets for MLLMs
Harnessing Instruction Tuning data can substantially bolster a model’s proficiency in adhering to instructions and enhancing its performance across specific tasks. Given the pivotal role of both quality and quantity in optimizing model performance, the exploration of this area emerges as a crucial research endeavor. The evaluation of data for large-scale models necessitates metrics that are model-agnostic, effectively capturing the distinct characteristics of the data. As highlighted in the preceding paragraph, there is a pressing need to devise metrics tailored specifically for instruction tuning data. An alternative pragmatic approach involves employing Large Language Models (LLMs) for data evaluation, leveraging the extensive knowledge accrued during their pre-training phase (Du et al., 2023; Chen et al., 2023f; Xu et al., 2023b; Liu et al., 2023e; Lu et al., 2023c; Parkar et al., 2024). However, utilizing LLMs for data assessment proves to be cost-effective. Despite the success of GPT-based methods for automated data quality evaluation, such approaches often lack interpretability. Hence, further investigation is warranted to gain a deeper understanding of GPT’s efficacy and to delineate the boundaries of employing GPT for data evaluation.
@@
MLLMs often learn from diverse data types sequentially across various training phases (Li et al., 2022a; Liu et al., 2024b, 2023c). Typically, this sequential training starts with initializing the model using pre-trained weights for each modality, followed by pre-training, instruction tuning, and concluding with reinforcement learning from human feedback. It is critical throughout these stages to prevent catastrophic forgetting, ensuring the model retains its foundational language capabilities while acquiring multimodal skills. Therefore, integrating lifelong learning strategies with MLLMs represents a significant and necessary research direction.